OpenStudy (anonymous):

Find an equation of the circle that satisfies: Center (7,-3); tangent to the x-axis

OpenStudy (anonymous):

hellooo?!

OpenStudy (anonymous):

PLEASE HELP ME!

OpenStudy (anonymous):

\[(x-7)^2 + (y+3)^2 = r^2\] That is the equation of a circle with center (7,-3) and radius r. Since the circle is tangent to the x-axis, that means the x-axis and the circle will touch at one point. Since the center is 3 units away from the x-axis, the radius will have to be 3 otherwise they would not be tangent. So the final equation is: \[(x-7)^2 + (y+3)^2 = 9\]